### v2

# Orbital-Selective Diffuse Magnetic Fluctuations in Sr2RuO4: a Unified Theoretical Picture

Abstract
The quasi-two-dimensional material Sr2RuO4 is a paradigmatic example of a correlated system that exhibits unconventional superconductivity and intriguing magnetic properties. The interplay between these two effects and the resulting strength and nature of spin fluctuations and their role for the properties of the compound have sparked significant debates. Here, elaborating a theory that self-consistently incorporates spatial magnetic fluctuations into a realistic many-body description, we show that these fluctuations significantly reduce many-body correlations in the system, thereby preventing magnetic ordering in Sr2RuO4, in agreement with experimental observations. Our conclusion is supported by a theoretical calculation of the spin susceptibility that closely matches the experimental results. We obtain finite peaks at the incommensurate wave vectors, a broad dome-shaped structure centered around the point and a diminished magnetic response at the edges of the BZ. We identify the orbital character of the unusual dome structure as resulting predominantly from the 2D-like -orbital, which is believed to be responsible for the superconductivity.
In materials with strong electronic correlations, typically containing partially filled electronic - or -shells, Coulomb interactions between the electrons play a dominant role in determining their properties. The behavior of electrons cannot be described independently from one another, as is mostly the case for conventional metals. Instead, complex collective phenomena emerge, including high-temperature superconductivity, Mott insulating states, or exotic magnetic phases.
The layered ruthenate Sr2RuO4 has attracted significant attention due to its structural similarity to high-temperature superconducting cuprates Keimer et al. (2015) and its unconventional superconductivity below K Maeno et al. (1994); Mackenzie and Maeno (2003); Maeno et al. (2012). Currently, ongoing efforts focus on understanding the enhancement of under uniaxial strain H. M. L. Noad and K. Ishida and Y.-S. Li and E. Gati and V. Stangier and N. Kikugawa and D. A. Sokolov and M. Nicklas and B. Kim and I. I. Mazin and M. Garst and J. Schmalian and A. P. Mackenzie and C. W. Hicks (2023); Maeno et al. (2024); Clifford W. Hicks and Daniel O. Brodsky and Edward A. Yelland and Alexandra S. Gibbs and Jan A. N. Bruin and Mark E. Barber and Stephen D. Edkins and Keigo Nishimura and Shingo Yonezawa and Yoshiteru Maeno and Andrew P. Mackenzie (2014); You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie (2021). Despite extensive research, however, the precise nature of the superconducting state in Sr2RuO4 remains unresolved. Conflicting experimental evidence suggests scenarios ranging from a single Chronister et al. (2021); Li et al. (2022); Jerzembeck et al. (2022); Palle et al. (2023) to broken time-reversal symmetry Xia et al. (2006); Luke et al. (1998), and a two-component Maeno et al. (2012); Kivelson et al. (2020); Maeno et al. (2024) order parameter. Numerous theoretical studies have tried to resolve this issue Rømer et al. (2019); Käser et al. (2022); Rømer et al. (2022); Profe et al. (2024), but no consensus has been reached so far.
At higher temperatures, significant magnetic fluctuations are found, and they are believed to be the source of the superconducting pairing mechanism in Sr2RuO4, calling urgently for their accurate theoretical description. As seen in inelastic neutron scattering (INS) measurements, the leading spin excitations in Sr2RuO4 correspond to the incommensurate wave vector Sidis et al. (1999); Braden et al. (2002a, 2004); Steffens et al. (2019); Jenni et al. (2021). Remarkably, although spin excitations in this material are exceptionally strong, different experimental studies confirm the absence of magnetic ordering in the parent compound. Nevertheless, Sr2RuO4 resides in close proximity to an ordered state, which can be induced already by a small concentration of impurities Minakata and Maeno (2001); Braden et al. (2002b); Kunkemöller et al. (2014), or by applying strain Grinenko et al. (2021); Li et al. (2022).
The description of spin excitations in Sr2RuO4 has sparked debates among theorists. This material is a paradigmatic correlated metal that requires an accurate non-perturbative treatment of electronic interactions. Conventional theoretical methods correctly capture the leading spin fluctuations associated with the incommensurate wave vector Mazin and Singh (1999); Boehnke et al. (2018); Acharya et al. (2019); Strand et al. (2019). However, they significantly overestimate the strength of magnetic fluctuations, predicting a transition to a magnetically ordered state that is not realized in nature. The final missing component in state-of-the-art theoretical calculations, which could potentially reconcile the discrepancies between theory and experiment, is a self-consistent incorporation of spatial magnetic fluctuations into the theory. This effect has not been addressed yet due to the lack of appropriate theoretical tools.
In this Letter, we resolve this long-standing problem by applying a recently developed advanced many-body approach that enables a self-consistent treatment of non-local collective electronic fluctuations within a realistic multi-orbital framework. We demonstrate that a self-consistent treatment of spatial spin excitations suppresses their strength and reduces the electronic correlations in Sr2RuO4, ultimately eliminating the magnetic ordering predicted by state-of-the-art approaches. Moreover, our results accurately reproduce the form of the spin susceptibility as deduced from experimental INS measurements, allowing us to detect an intriguing orbital dependence of the latter.
The Fermi surface (FS) of Sr2RuO4, shown in Fig. 1, consists of three Fermi sheets, originating from the three orbitals: the quasi-one-dimensional (1D) and , and the quasi-two-dimensional (2D) Mackenzie et al. (1996); Bergemann et al. (2000); Damascelli et al. (2000); Tamai et al. (2019). Angle-resolved photoemission spectroscopy (ARPES) and quantum oscillations measurements reveal that Sr2RuO4 is a correlated metal Ingle et al. (2005); Damascelli et al. (2000); Mackenzie et al. (1996). Although a significant redistribution of the spectral weight and an enhancement of it around the Fermi energy are observed Lu et al. (1996); Yokoya et al. (1996), the band structure and FS remain relatively close Mazin and Singh (1997); Pavarini and Mazin (2006) to those predicted by density functional theory (DFT) Kohn and Sham (1965). From the theoretical perspective, the compound is considered to be a paradigmatic correlated Fermi liquid, with strong signatures of Hund’s metal physics Mravlje et al. (2011); Georges et al. (2013); Kugler et al. (2020); Hunter et al. (2023); Blesio, Germán and Beck, Sophie and Gingras, Olivier and Georges, Antoine and Mravlje, Jernej (2024); Suzuki et al. (2023). Therefore, a number of studies have been conducted using this material as a testbed for realistic many-body calculations.
Early electronic structure calculations using DFT combined with dynamical mean-field theory (DMFT) Georges et al. (1996), that accounts for local correlation effects, have nearly succeeded to reproduce the experimental spectral function Liebsch and Lichtenstein (2000). Since then, several DMFT studies have allowed for many important advances in our understanding of the physics of Sr2RuO4 Mravlje et al. (2011); Hunter et al. (2023); Zhang et al. (2016). Recent studies Tamai et al. (2019); Hunter et al. (2023) showed that the self-energy, derived from photoemission measurements at the k-points of the Brillouin Zone (BZ) corresponding to the FS, is predominantly local for the orbitals. In contrast, the orbital exhibits some momentum dependence, which becomes apparent at frequencies above meV Tamai et al. (2019). These findings suggest that, overall, the compound’s single-particle properties are not particularly remarkable and are adequately captured by DMFT.
It is therefore very surprising that the material’s two-particle quantities, particularly spin excitations, are in contrast not well understood. Within DFT calculations, the FS nesting arising from the compound’s crystal structure leads to an ordered spin-density wave (SDW) state with the incommensurate wave vectors Mazin and Singh (1999). DMFT calculations of the spin susceptibility, incorporating local vertex corrections, have also reproduced the peaks at the vectors Boehnke et al. (2018); Acharya et al. (2019); Strand et al. (2019). However, DMFT strongly overestimates the strength of magnetic fluctuations, compared to experiments, predicting a transition to a SDW ordered state at a finite temperature K Strand et al. (2019). While it has been shown that spin-orbit coupling (SOC) can suppress this magnetic transition in DMFT, achieving this requires a rather large coupling strength Strand et al. (2019). Additionally, the SOC is not expected to significantly influence the spin susceptibility at the relatively high temperatures at which the DMFT calculations have been conducted Eremin et al. (2002); Braden et al. (2004).
The key contribution of DMFT to calculating the spin susceptibility is its ability to reveal the suppression of the magnetic signal at the edges of the BZ, particularly at the point, in agreement with experiments. Within DMFT, spin excitations, aside from the peaks, are found to be quasi-local, or nearly constant in momentum space. However, this result does not fully align with experimental findings, where the magnetic response, in addition to the peaks, exhibits a relatively broad dome centered at the point Steffens et al. (2019); Jenni et al. (2021, 2021), rather than a quasi-local background signal.
To accurately describe magnetic fluctuations in Sr2RuO4, we consider an effective three-band model (see the Supplemental Material (SM) SM ) corresponding to maximally localized orbitals derived from DFT 111 The DFT calculations were performed within the projected augmented wave formalism Blöchl (1994); Kresse and Joubert (1999) as implemented in the Vienna ab initio simulation package (VASP) Kresse and Furthmüller (1996, 1996). The exchange-correlation effects were considered within the generalized-gradient approximation functional in the PBE parametrization Perdew et al. (1996). We used the standard pseudopotentials that include 10, 16, and 6 valence electrons for Sr, Ru, and O, respectively. An energy cut-off of 400 eV for the plane-waves and a convergence threshold of eV were used in the calculations. The Wannier functions and the effective tight-binding Hamiltonian were constructed within the scheme of maximal localization Marzari and Vanderbilt (1997); Marzari et al. (2012) using the wannier90 package Mostofi et al. (2008). . We account for the on-site electronic interaction that is parametrized in the Kanamori form Kanamori (1963). The value of the intra-orbital Coulomb repulsion, eV, is chosen based on the constrained random-phase approximation (cRPA) analysis conducted in Vaugier et al. (2012). A Hund’s exchange coupling , crucial for Sr2RuO4, is selected by evaluating results across different values of . Specifically, we estimate the mass enhancement and spin susceptibility obtained for each , as detailed in the SM SM . We find that the correct mass enhancement is achieved for eV, and the accurate spin susceptibility for eV, leading to an optimal choice of eV. We do not include spin-orbit coupling (SOC) in our calculations because it is prohibitively expensive computationally. This approximation is supported by experimental evidences, suggesting that SOC effects become less significant at higher temperatures, due to thermal effects.
In order to obtain a reliable description of the feedback of electronic correlations on the spectral and magnetic properties of the compound, a self-consistent method is required. We employ the dual triply irreducible local expansion (D-TRILEX) approach Stepanov et al. (2019); Harkov et al. (2021); Vandelli et al. (2022), which is a diagrammatic extension of DMFT Rohringer et al. (2018); Lyakhova et al. (2023). The decisive advantage of this method over other DMFT extensions is its computational efficiency, which enables calculations for multi-orbital systems Stepanov et al. (2021); Vandelli et al. (2023); Stepanov (2022); Stepanov and Biermann (2024); Stepanov et al. (2024a), such as Sr2RuO4. In the D-TRILEX framework, local electronic correlations are treated non-perturbatively via a DMFT impurity problem Georges et al. (1996), which is solved using the w2dynamics package Wallerberger et al. (2019). The latter further serves as a reference system for the diagrammatic expansion that self-consistently incorporates non-local electronic correlations, including magnetic fluctuations. This approach allows the mutual influence of collective electronic fluctuations on single-particle quantities and vice versa Stepanov et al. (2022); Vandelli et al. (2024); Chatzieleftheriou et al. (2024); Stepanov et al. (2024b), yielding reliable results for both single- and two-particle observables in a broad regime of system’s parameters Harkov et al. (2021); Vandelli et al. (2022); Vandelli (2022).
We perform the D-TRILEX calculations at a relatively low temperature K, where experimental data are available for comparison. The inset in Fig. 2 shows the real part of the static magnetic () susceptibility obtained for the first BZ using D-TRILEX. The main part of Fig. 2 displays a cut of the susceptibility along the X--X diagonal of the BZ (dashed line in the inset). The results are calculated numerically using D-TRILEX (blue) and DMFT (green), and are compared to the INS result (magenta) Steffens et al. (2019). In the INS study, performed at K, the real part of the static spin susceptibility is deduced from the fit of the low energy part of the experimentally measured spin excitation spectrum, with a set of single relaxors as described in Ref. Steffens et al. (2019). At this temperature the magnetic fluctuations are already very strong. In fact, for the considered model DMFT predicts a SDW ordered state already at K 222To be precise, our DMFT susceptibility is obtained through a single-shot D-TRILEX calculation, which, in practice, gives a very similar result to a direct evaluation of the local vertex corrections, as done in Ref. Strand et al. (2019). In that work the transition temperature is found to be lower at , but this is related to the different interaction parameters used in the current study.. For this reason, the DMFT result in Fig. 2 is shown for a bit higher temperature K. The spin susceptibility of DMFT, calculated in the vicinity of the SDW transition, features large peaks at the incommensurate vectors, in agreement with previous works Strand et al. (2019). For easier comparison, the DMFT result in Fig. 2 is multiplied by . We find, that a self-consistent inclusion of the magnetic fluctuations beyond DMFT, using D-TRILEX, leads to a strong suppression of the SDW peaks in a good agreement with the INS result, and no ordering is observed.
The second important outcome of our results is related to the overall behavior of the spin susceptibility across the BZ. As discussed above, according to experimental measurements the magnetic signal can be decomposed into the sum of the SDW peaks and a broad dome structure centered around the point, while DMFT calculations instead find a quasi-constant background signal besides the peaks. The D-TRILEX calculations reveal a significantly diminished spin susceptibility at the edges of the BZ, with a “cross”-like structure in momentum space of higher intensity appearing at the center of the zone, visible in the inset of Fig. 2. The overall structure of the susceptibility agrees very well with the experimental results Steffens et al. (2019). While early-unpolarized INS studies reported features of the spin excitation spectrum around Braden et al. (2002a), the most recent polarized INS data Steffens et al. (2019); Jenni et al. (2021) unambiguously confirmed the existence of such fluctuations, but cannot resolve the actual structure around . Nevertheless, strong spin excitations around are found in metamagnetic Ca2-xSrxRuO4 Friedt et al. (2004); Steffens et al. (2007, 2011), where the isostructural Ca/Sr substitution is, however, accompanied by a rotation and deformation of the RuO6 octahedra Friedt et al. (2001); Kriener et al. (2005).
In Fig. 3 we show the orbitally resolved static spin susceptibility , where . The results are calculated using D-TRILEX (blue) and DMFT (green) and plotted along the X--X path. Although the incommensurate SDW peaks are present in all orbital components due to mixing of different of orbital contributions to the susceptibility, the value of is the largest for the intra-orbital - and - components (a). This is expected, since the SDW peaks originate from the nesting of the FS of these 1D-like orbitals. Interestingly, the background dome-shaped magnetic signal, found experimentally and reproduced by D-TRILEX, appears to originate from magnetic fluctuations within the 2D-like orbital (b). Instead, the DMFT results do not exhibit any significant signal around the point, and for all the components a quasi-local response is obtained besides the peaks. This result may have significant implications for the symmetry of the superconducting order parameter, as recent studies suggest that magnetic fluctuations within the orbital are a key ingredient for the electron pairing Rømer et al. (2019); Profe et al. (2024). In particular, the presence of spin excitations at more than one wave-vector allows for a competition between different order parameters of the superconducting state Ghosh et al. (2021); Benhabib et al. (2021).
The breakdown of the SDW ordering predicted by DMFT, upon self-consistent inclusion of magnetic fluctuations in D-TRILEX, can be understood as follows. In a metal, strong spin fluctuations induce a large electronic self-energy, which in turn diminishes the electronic Green’s function. This renders the system less correlated, thereby decreasing the electronic polarizability, which renormalizes the spin excitations and thus suppresses them. The reduction of the many-body effects by magnetic fluctuations in Sr2RuO4 is clearly reflected in its single-particle properties. In Fig. 1 (e, f) we plot the imaginary part of the electronic Green’s function at the zeroth Matsubara frequency , which approximates the FS. The result is calculated using D-TRILEX (e) and DMFT (f) at K. Both methods reveal three FS sheets originating from the 1D / and the 2D orbitals. However, one immediately finds that DMFT predicts a large broadening of the FS, while D-TRILEX shows instead very sharp FS sheets. In Fig. 4 we plot the momentum-resolved spectral function along the high-symmetry path -M-X- in the BZ. One immediately notes that DMFT (b) significantly renormalizes the non-interacting DFT band structure (white lines), particularly the bandwidth, and pins the van Hove singularity to the Fermi level. The latter is a well known effect of strong electronic correlations seen in various systems Liechtenstein et al. (1996); Irkhin et al. (2002); Odashima et al. (2005); Chen et al. (2011); Schmitt (2010); Stepanov et al. (2024a). The renormalization predicted by DMFT broadens the bands and increases the density of states around the Fermi energy, strongly enhancing the leading nesting-driven spin excitations and ultimately resulting in a phase transition to a magnetically ordered state. Essentially, DMFT appears to overestimate the degree of correlation of the system. Instead, the self-consistent inclusion of spatial magnetic fluctuations within D-TRILEX reduces the electronic correlations, resulting in sharper, less renormalized bands (a) that are rather close to the DFT picture. In particular, we find that the van Hove singularity is no longer pinned to the Fermi level but appears above the Fermi energy, as in DFT, indicating that electronic correlations within the D-TRILEX framework are weaker than those in DMFT.
The similarity between the D-TRILEX and DFT band structures also explains the emergence of the dome-shaped -point structure in the spin susceptibility of D-TRILEX, as a similar signal has been observed in RPA calculations for the DFT band structure under specific conditions Rømer et al. (2019). The proximity of the spectral function of Sr2RuO4 to the DFT one is also reflected in the orbital polarization. The latter is found to be decreased in DMFT, yielding the values , compared to the DFT values . This significant polarization between the 1D and 2D orbitals is also observed in D-TRILEX (), and it is another signature of reduced electronic correlations.
By analyzing the electron self-energy we find that, although it is not local and exhibits a noticeable momentum dependence perpendicularly to the FS, it has a surprisingly small momentum-dependence along the FS, as discussed in the SM SM . Along the FS the / orbitals appear to have a practically constant self-energy, while the orbital exhibits a finite but small momentum-modulation. This result is in a qualitative agreement with recent ARPES data presented in Fig. 4 (d) of Ref. Tamai et al. (2019), and explains why DMFT was particularly successful in reproducing the FS of the material.
In conclusion, we have studied the effect of magnetic fluctuations on the electronic correlations and the spin susceptibility of Sr2RuO4.
These excitations are found to be significant in this material, such that DMFT calculations predict an ordered SDW state that is not observed experimentally.
We demonstrate that the self-consistent inclusion of spatial magnetic fluctuations suppresses their strength by reducing many-body correlations to the extent that no magnetic ordering is realized in Sr2RuO4.
The overall behavior of the spin susceptibility in momentum space deduced from INS measurements is well reproduced by our calculations.
We obtain finite peaks at the incommensurate SDW vectors, a broad dome-shaped structure centered around the point and a diminished magnetic response at the edges of the BZ.
We identify the orbital character of the unusual dome structure as resulting predominantly from the 2D-like -orbital. This observation resonates with recent claims that it is the -orbital that
is mainly responsible for the superconductivity.
Acknowledgements.
We thank Antoine Georges, Alexander Lichtenstein, Hugo Strand and Alaska Subedi for fruitful discussions. We acknowledge support from IDRIS/GENCI under grant number 091393, and are thankful to the CPHT computer support team for their help.References
- Keimer et al. (2015) B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, “From quantum matter to high-temperature superconductivity in copper oxides,” Nature 518, 179–186 (2015).
- Maeno et al. (1994) Y. Maeno, H. Hashimoto, K. Yoshida, S. Nishizaki, T. Fujita, J. G. Bednorz, and F. Lichtenberg, “Superconductivity in a layered perovskite without copper,” Nature 372, 532–534 (1994).
- Mackenzie and Maeno (2003) Andrew Peter Mackenzie and Yoshiteru Maeno, “The superconductivity of and the physics of spin-triplet pairing,” Rev. Mod. Phys. 75, 657–712 (2003).
- Maeno et al. (2012) Yoshiteru Maeno, Shunichiro Kittaka, Takuji Nomura, Shingo Yonezawa, and Kenji Ishida, “Evaluation of Spin-Triplet Superconductivity in Sr2RuO4,” J. Phys. Soc. Jpn. 81, 011009 (2012).
- H. M. L. Noad and K. Ishida and Y.-S. Li and E. Gati and V. Stangier and N. Kikugawa and D. A. Sokolov and M. Nicklas and B. Kim and I. I. Mazin and M. Garst and J. Schmalian and A. P. Mackenzie and C. W. Hicks (2023) H. M. L. Noad and K. Ishida and Y.-S. Li and E. Gati and V. Stangier and N. Kikugawa and D. A. Sokolov and M. Nicklas and B. Kim and I. I. Mazin and M. Garst and J. Schmalian and A. P. Mackenzie and C. W. Hicks , “Giant lattice softening at a Lifshitz transition in Sr2RuO4,” Science 382, 447–450 (2023).
- Maeno et al. (2024) Yoshiteru Maeno, Shingo Yonezawa, and Aline Ramires, “Still Mystery after All These Years —Unconventional Superconductivity of Sr2RuO4—,” J. Phys. Soc. Jpn. 93, 062001 (2024).
- Clifford W. Hicks and Daniel O. Brodsky and Edward A. Yelland and Alexandra S. Gibbs and Jan A. N. Bruin and Mark E. Barber and Stephen D. Edkins and Keigo Nishimura and Shingo Yonezawa and Yoshiteru Maeno and Andrew P. Mackenzie (2014) Clifford W. Hicks and Daniel O. Brodsky and Edward A. Yelland and Alexandra S. Gibbs and Jan A. N. Bruin and Mark E. Barber and Stephen D. Edkins and Keigo Nishimura and Shingo Yonezawa and Yoshiteru Maeno and Andrew P. Mackenzie , “Strong Increase of of Sr2RuO4 Under Both Tensile and Compressive Strain,” Science 344, 283–285 (2014).
- You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie (2021) You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie , “High-sensitivity heat-capacity measurements on Sr2RuO4 under uniaxial pressure,” Proc. Natl. Acad. Sci. U.S.A. 118, e2020492118 (2021).
- Chronister et al. (2021) Aaron Chronister, Andrej Pustogow, Naoki Kikugawa, Dmitry A. Sokolov, Fabian Jerzembeck, Clifford W. Hicks, Andrew P. Mackenzie, Eric D. Bauer, and Stuart E. Brown, “Evidence for even parity unconventional superconductivity in Sr2RuO4,” Proc. Natl. Acad. Sci. U.S.A. 118, e2025313118 (2021).
- Li et al. (2022) You-Sheng Li, Markus Garst, Jörg Schmalian, Sayak Ghosh, Naoki Kikugawa, Dmitry A Sokolov, Clifford W Hicks, Fabian Jerzembeck, Matthias S Ikeda, Zhenhai Hu, et al., “Elastocaloric determination of the phase diagram of Sr2RuO4,” Nature 607, 276–280 (2022).
- Jerzembeck et al. (2022) Fabian Jerzembeck, Henrik S. Røising, Alexander Steppke, Helge Rosner, Dmitry A. Sokolov, Naoki Kikugawa, Thomas Scaffidi, Steven H. Simon, Andrew P. Mackenzie, and Clifford W. Hicks, “The superconductivity of Sr2RuO4 under c-axis uniaxial stress,” Nat. Commun. 13, 4596 (2022).
- Palle et al. (2023) Grgur Palle, Clifford Hicks, Roser Valentí, Zhenhai Hu, You-Sheng Li, Andreas Rost, Michael Nicklas, Andrew P. Mackenzie, and Jörg Schmalian, “Constraints on the superconducting state of from elastocaloric measurements,” Phys. Rev. B 108, 094516 (2023).
- Xia et al. (2006) Jing Xia, Yoshiteru Maeno, Peter T. Beyersdorf, M. M. Fejer, and Aharon Kapitulnik, “High Resolution Polar Kerr Effect Measurements of : Evidence for Broken Time-Reversal Symmetry in the Superconducting State,” Phys. Rev. Lett. 97, 167002 (2006).
- Luke et al. (1998) G. Ml. Luke, Y. Fudamoto, K. M. Kojima, M. I. Larkin, J. Merrin, B. Nachumi, Y. J. Uemura, Y. Maeno, Z. Q. Mao, Y. Mori, H. Nakamura, and M. Sigrist, “Time-reversal symmetry-breaking superconductivity in Sr2RuO4,” Nature 394, 558–561 (1998).
- Kivelson et al. (2020) Steven Allan Kivelson, Andrew Chang Yuan, Brad Ramshaw, and Ronny Thomale, “A proposal for reconciling diverse experiments on the superconducting state in Sr2RuO4,” npj Quantum Mater. 5, 43 (2020).
- Rømer et al. (2019) A. T. Rømer, D. D. Scherer, I. M. Eremin, P. J. Hirschfeld, and B. M. Andersen, “Knight Shift and Leading Superconducting Instability from Spin Fluctuations in ,” Phys. Rev. Lett. 123, 247001 (2019).
- Käser et al. (2022) Stefan Käser, Hugo U. R. Strand, Nils Wentzell, Antoine Georges, Olivier Parcollet, and Philipp Hansmann, “Interorbital singlet pairing in : A Hund’s superconductor,” Phys. Rev. B 105, 155101 (2022).
- Rømer et al. (2022) Astrid T. Rømer, T. A. Maier, Andreas Kreisel, P. J. Hirschfeld, and Brian M. Andersen, “Leading superconducting instabilities in three-dimensional models for ,” Phys. Rev. Res. 4, 033011 (2022).
- Profe et al. (2024) Jonas B. Profe, Sophie Beck, Dante M. Kennes, Antoine Georges, and Olivier Gingras, “Competition between d-wave superconductivity and magnetism in uniaxially strained Sr2RuO4,” npj Quantum Mater. 9, 53 (2024).
- Sidis et al. (1999) Y. Sidis, M. Braden, P. Bourges, B. Hennion, S. NishiZaki, Y. Maeno, and Y. Mori, “Evidence for Incommensurate Spin Fluctuations in ,” Phys. Rev. Lett. 83, 3320–3323 (1999).
- Braden et al. (2002a) M. Braden, Y. Sidis, P. Bourges, P. Pfeuty, J. Kulda, Z. Mao, and Y. Maeno, “Inelastic neutron scattering study of magnetic excitations in ,” Phys. Rev. B 66, 064522 (2002a).
- Braden et al. (2004) M. Braden, P. Steffens, Y. Sidis, J. Kulda, P. Bourges, S. Hayden, N. Kikugawa, and Y. Maeno, “Anisotropy of the Incommensurate Fluctuations in : A Study with Polarized Neutrons,” Phys. Rev. Lett. 92, 097402 (2004).
- Steffens et al. (2019) P. Steffens, Y. Sidis, J. Kulda, Z. Q. Mao, Y. Maeno, I. I. Mazin, and M. Braden, “Spin Fluctuations in from Polarized Neutron Scattering: Implications for Superconductivity,” Phys. Rev. Lett. 122, 047004 (2019).
- Jenni et al. (2021) K. Jenni, S. Kunkemöller, P. Steffens, Y. Sidis, R. Bewley, Z. Q. Mao, Y. Maeno, and M. Braden, “Neutron scattering studies on spin fluctuations in ,” Phys. Rev. B 103, 104511 (2021).
- Minakata and Maeno (2001) M. Minakata and Y. Maeno, “Magnetic ordering in induced by nonmagnetic impurities,” Phys. Rev. B 63, 180504 (2001).
- Braden et al. (2002b) M. Braden, O. Friedt, Y. Sidis, P. Bourges, M. Minakata, and Y. Maeno, “Incommensurate Magnetic Ordering in ,” Phys. Rev. Lett. 88, 197002 (2002b).
- Kunkemöller et al. (2014) S. Kunkemöller, A. A. Nugroho, Y. Sidis, and M. Braden, “Spin-density-wave ordering in Sr1.5 studied by neutron scattering,” Phys. Rev. B 89, 045119 (2014).
- Grinenko et al. (2021) Vadim Grinenko, Shreenanda Ghosh, Rajib Sarkar, Jean-Christophe Orain, Artem Nikitin, Matthias Elender, Debarchan Das, Zurab Guguchia, Felix Brückner, Mark E Barber, et al., “Split superconducting and time-reversal symmetry-breaking transitions in Sr2RuO4 under stress,” Nat. Phys. 17, 748–754 (2021).
- Mazin and Singh (1999) I. I. Mazin and D. J. Singh, “Competitions in layered ruthenates: Ferromagnetism versus antiferromagnetism and triplet versus singlet pairing,” Phys. Rev. Lett. 82, 4324–4327 (1999).
- Boehnke et al. (2018) Lewin Boehnke, Philipp Werner, and Frank Lechermann, “Multi-orbital nature of the spin fluctuations in Sr2RuO4,” EPL 122, 57001 (2018).
- Acharya et al. (2019) Swagata Acharya, Dimitar Pashov, Cédric Weber, Hyowon Park, Lorenzo Sponza, and Mark Van Schilfgaarde, “Evening out the spin and charge parity to increase in Sr2RuO4,” Commun. Phys. 2, 163 (2019).
- Strand et al. (2019) Hugo U. R. Strand, Manuel Zingl, Nils Wentzell, Olivier Parcollet, and Antoine Georges, “Magnetic response of : Quasi-local spin fluctuations due to Hund’s coupling,” Phys. Rev. B 100, 125120 (2019).
- Mackenzie et al. (1996) A. P. Mackenzie, S. R. Julian, A. J. Diver, G. J. McMullan, M. P. Ray, G. G. Lonzarich, Y. Maeno, S. Nishizaki, and T. Fujita, “Quantum Oscillations in the Layered Perovskite Superconductor SRu,” Phys. Rev. Lett. 76, 3786–3789 (1996).
- Bergemann et al. (2000) C. Bergemann, S. R. Julian, A. P. Mackenzie, S. NishiZaki, and Y. Maeno, “Detailed Topography of the Fermi Surface of ,” Phys. Rev. Lett. 84, 2662–2665 (2000).
- Damascelli et al. (2000) A. Damascelli, D. H. Lu, K. M. Shen, N. P. Armitage, F. Ronning, D. L. Feng, C. Kim, Z.-X. Shen, T. Kimura, Y. Tokura, Z. Q. Mao, and Y. Maeno, “Fermi Surface, Surface States, and Surface Reconstruction in ,” Phys. Rev. Lett. 85, 5194–5197 (2000).
- Tamai et al. (2019) A. Tamai, M. Zingl, E. Rozbicki, E. Cappelli, S. Riccò, A. de la Torre, S. McKeown Walker, F. Y. Bruno, P. D. C. King, W. Meevasana, M. Shi, M. Radović, N. C. Plumb, A. S. Gibbs, A. P. Mackenzie, C. Berthod, H. U. R. Strand, M. Kim, A. Georges, and F. Baumberger, “High-Resolution Photoemission on Reveals Correlation-Enhanced Effective Spin-Orbit Coupling and Dominantly Local Self-Energies,” Phys. Rev. X 9, 021048 (2019).
- Ingle et al. (2005) N. J. C. Ingle, K. M. Shen, F. Baumberger, W. Meevasana, D. H. Lu, Z.-X. Shen, A. Damascelli, S. Nakatsuji, Z. Q. Mao, Y. Maeno, T. Kimura, and Y. Tokura, “Quantitative analysis of angle-resolved photoemission spectra: Many-body interactions in a model Fermi liquid,” Phys. Rev. B 72, 205114 (2005).
- Lu et al. (1996) D. H. Lu, M. Schmidt, T. R. Cummins, S. Schuppler, F. Lichtenberg, and J. G. Bednorz, “Fermi Surface and Extended van Hove Singularity in the Noncuprate Superconductor SRu,” Phys. Rev. Lett. 76, 4845–4848 (1996).
- Yokoya et al. (1996) T. Yokoya, A. Chainani, T. Takahashi, H. Katayama-Yoshida, M. Kasai, and Y. Tokura, “Extended Van Hove Singularity in a Noncuprate Layered Superconductor SRu,” Phys. Rev. Lett. 76, 3009–3012 (1996).
- Mazin and Singh (1997) I. I. Mazin and David J. Singh, “Ferromagnetic Spin Fluctuation Induced Superconductivity in ,” Phys. Rev. Lett. 79, 733–736 (1997).
- Pavarini and Mazin (2006) E. Pavarini and I. I. Mazin, “First-principles study of spin-orbit effects and NMR in ,” Phys. Rev. B 74, 035115 (2006).
- Kohn and Sham (1965) W. Kohn and L. J. Sham, “Self-consistent equations including exchange and correlation effects,” Phys. Rev. 140, A1133–A1138 (1965).
- Mravlje et al. (2011) Jernej Mravlje, Markus Aichhorn, Takashi Miyake, Kristjan Haule, Gabriel Kotliar, and Antoine Georges, “Coherence-Incoherence Crossover and the Mass-Renormalization Puzzles in ,” Phys. Rev. Lett. 106, 096401 (2011).
- Georges et al. (2013) Antoine Georges, Luca de’ Medici, and Jernej Mravlje, “Strong Correlations from Hund’s Coupling,” Annu. Rev. Condens. Matter Phys. 4, 137–178 (2013).
- Kugler et al. (2020) Fabian B. Kugler, Manuel Zingl, Hugo U. R. Strand, Seung-Sup B. Lee, Jan von Delft, and Antoine Georges, “Strongly Correlated Materials from a Numerical Renormalization Group Perspective: How the Fermi-Liquid State of Emerges,” Phys. Rev. Lett. 124, 016401 (2020).
- Hunter et al. (2023) A. Hunter, S. Beck, E. Cappelli, F. Margot, M. Straub, Y. Alexanian, G. Gatti, M. D. Watson, T. K. Kim, C. Cacho, N. C. Plumb, M. Shi, M. Radović, D. A. Sokolov, A. P. Mackenzie, M. Zingl, J. Mravlje, A. Georges, F. Baumberger, and A. Tamai, “Fate of Quasiparticles at High Temperature in the Correlated Metal Sr2RuO4,” Phys. Rev. Lett. 131, 236502 (2023).
- Blesio, Germán and Beck, Sophie and Gingras, Olivier and Georges, Antoine and Mravlje, Jernej (2024) Blesio, Germán and Beck, Sophie and Gingras, Olivier and Georges, Antoine and Mravlje, Jernej, “Signatures of Hund metal and finite-frequency nesting in revealed by electronic Raman scattering,” Phys. Rev. Res. 6, 023124 (2024).
- Suzuki et al. (2023) H. Suzuki, L. Wang, J. Bertinshaw, H. U. R. Strand, S. Käser, M. Krautloher, Z. Yang, N. Wentzell, O. Parcollet, F. Jerzembeck, N. Kikugawa, A. P. Mackenzie, A. Georges, P. Hansmann, H. Gretarsson, and B. Keimer, “Distinct spin and orbital dynamics in Sr2RuO4,” Nat. Commun. 14, 7042 (2023).
- Georges et al. (1996) Antoine Georges, Gabriel Kotliar, Werner Krauth, and Marcelo J. Rozenberg, “Dynamical mean-field theory of strongly correlated fermion systems and the limit of infinite dimensions,” Rev. Mod. Phys. 68, 13–125 (1996).
- Liebsch and Lichtenstein (2000) A. Liebsch and A. Lichtenstein, “Photoemission Quasiparticle Spectra of Sr2RuO4,” Phys. Rev. Lett. 84, 1591–1594 (2000).
- Zhang et al. (2016) Guoren Zhang, Evgeny Gorelov, Esmaeel Sarvestani, and Eva Pavarini, “Fermi Surface of : Spin-Orbit and Anisotropic Coulomb Interaction Effects,” Phys. Rev. Lett. 116, 106402 (2016).
- Eremin et al. (2002) I. Eremin, D. Manske, and K. H. Bennemann, “Electronic theory for the normal-state spin dynamics in Anisotropy due to spin-orbit coupling,” Phys. Rev. B 65, 220502 (2002).
- (53) See Supplemental Material for an analysis of the effect of Hund’s exchange coupling and of temperature on the system.
- Note (1) The DFT calculations were performed within the projected augmented wave formalism Blöchl (1994); Kresse and Joubert (1999) as implemented in the Vienna ab initio simulation package (VASP) Kresse and Furthmüller (1996, 1996). The exchange-correlation effects were considered within the generalized-gradient approximation functional in the PBE parametrization Perdew et al. (1996). We used the standard pseudopotentials that include 10, 16, and 6 valence electrons for Sr, Ru, and O, respectively. An energy cut-off of 400 eV for the plane-waves and a convergence threshold of eV were used in the calculations. The Wannier functions and the effective tight-binding Hamiltonian were constructed within the scheme of maximal localization Marzari and Vanderbilt (1997); Marzari et al. (2012) using the wannier90 package Mostofi et al. (2008).
- Kanamori (1963) Junjiro Kanamori, “Electron Correlation and Ferromagnetism of Transition Metals,” Prog. Theor. Phys. 30, 275–289 (1963).
- Vaugier et al. (2012) Loïg Vaugier, Hong Jiang, and Silke Biermann, “Hubbard and Hund exchange in transition metal oxides: Screening versus localization trends from constrained random phase approximation,” Phys. Rev. B 86, 165105 (2012).
- Stepanov et al. (2019) E. A. Stepanov, V. Harkov, and A. I. Lichtenstein, “Consistent partial bosonization of the extended Hubbard model,” Phys. Rev. B 100, 205115 (2019).
- Harkov et al. (2021) V. Harkov, M. Vandelli, S. Brener, A. I. Lichtenstein, and E. A. Stepanov, “Impact of partially bosonized collective fluctuations on electronic degrees of freedom,” Phys. Rev. B 103, 245123 (2021).
- Vandelli et al. (2022) Matteo Vandelli, Josef Kaufmann, Mohammed El-Nabulsi, Viktor Harkov, Alexander I. Lichtenstein, and Evgeny A. Stepanov, “Multi-band D-TRILEX approach to materials with strong electronic correlations,” SciPost Phys. 13, 036 (2022).
- Rohringer et al. (2018) G. Rohringer, H. Hafermann, A. Toschi, A. A. Katanin, A. E. Antipov, M. I. Katsnelson, A. I. Lichtenstein, A. N. Rubtsov, and K. Held, “Diagrammatic routes to nonlocal correlations beyond dynamical mean field theory,” Rev. Mod. Phys. 90, 025003 (2018).
- Lyakhova et al. (2023) Ya. S. Lyakhova, G. V. Astretsov, and A. N. Rubtsov, “The mean-field concept and post-DMFT methods in the contemporary theory of correlated systems,” Phys. Usp. 66, 775–793 (2023).
- Stepanov et al. (2021) Evgeny A. Stepanov, Yusuke Nomura, Alexander I. Lichtenstein, and Silke Biermann, “Orbital Isotropy of Magnetic Fluctuations in Correlated Electron Materials Induced by Hund’s Exchange Coupling,” Phys. Rev. Lett. 127, 207205 (2021).
- Vandelli et al. (2023) M. Vandelli, J. Kaufmann, V. Harkov, A. I. Lichtenstein, K. Held, and E. A. Stepanov, “Extended regime of metastable metallic and insulating phases in a two-orbital electronic system,” Phys. Rev. Res. 5, L022016 (2023).
- Stepanov (2022) Evgeny A. Stepanov, “Eliminating Orbital Selectivity from the Metal-Insulator Transition by Strong Magnetic Fluctuations,” Phys. Rev. Lett. 129, 096404 (2022).
- Stepanov and Biermann (2024) Evgeny A. Stepanov and Silke Biermann, “Can Orbital-Selective Néel Transitions Survive Strong Nonlocal Electronic Correlations?” Phys. Rev. Lett. 132, 226501 (2024).
- Stepanov et al. (2024a) Evgeny A. Stepanov, Matteo Vandelli, Alexander I. Lichtenstein, and Frank Lechermann, “Charge Density Wave Ordering in NdNiO2: Effects of Multiorbital Nonlocal Correlations,” npj Comput. Mater. 10, 108 (2024a).
- Wallerberger et al. (2019) Markus Wallerberger, Andreas Hausoel, Patrik Gunacker, Alexander Kowalski, Nicolaus Parragh, Florian Goth, Karsten Held, and Giorgio Sangiovanni, “w2dynamics: Local one- and two-particle quantities from dynamical mean field theory,” Comput. Phys. Commun. 235, 388–399 (2019).
- Stepanov et al. (2022) E. A. Stepanov, V. Harkov, M. Rösner, A. I. Lichtenstein, M. I. Katsnelson, and A. N. Rudenko, “Coexisting charge density wave and ferromagnetic instabilities in monolayer InSe,” npj Comput. Mater. 8, 118 (2022).
- Vandelli et al. (2024) M. Vandelli, A. Galler, A. Rubio, A. I. Lichtenstein, S. Biermann, and E. A. Stepanov, “Doping-dependent charge- and spin-density wave orderings in a monolayer of Pb adatoms on Si(111),” npj Quantum Mater. 9, 19 (2024).
- Chatzieleftheriou et al. (2024) Maria Chatzieleftheriou, Silke Biermann, and Evgeny A. Stepanov, “Local and Nonlocal Electronic Correlations at the Metal-Insulator Transition in the Two-Dimensional Hubbard Model,” Phys. Rev. Lett. 132, 236504 (2024).
- Stepanov et al. (2024b) Evgeny A. Stepanov, Maria Chatzieleftheriou, Niklas Wagner, and Giorgio Sangiovanni, “Interconnected renormalization of Hubbard bands and Green’s function zeros in Mott insulators induced by strong magnetic fluctuations,” Phys. Rev. B 110, L161106 (2024b).
- Vandelli (2022) Matteo Vandelli, Quantum embedding methods in dual space for strongly interacting electronic systems, Ph.D. thesis, Universität Hamburg Hamburg (2022).
- Note (2) To be precise, our DMFT susceptibility is obtained through a single-shot D-TRILEX calculation, which, in practice, gives a very similar result to a direct evaluation of the local vertex corrections, as done in Ref. Strand et al. (2019). In that work the transition temperature is found to be lower at , but this is related to the different interaction parameters used in the current study.
- Friedt et al. (2004) O. Friedt, P. Steffens, M. Braden, Y. Sidis, S. Nakatsuji, and Y. Maeno, “Strongly Enhanced Magnetic Fluctuations in a Large-Mass Layered Ruthenate,” Phys. Rev. Lett. 93, 147404 (2004).
- Steffens et al. (2007) P. Steffens, Y. Sidis, P. Link, K. Schmalzl, S. Nakatsuji, Y. Maeno, and M. Braden, “Field-Induced Paramagnons at the Metamagnetic Transition of ,” Phys. Rev. Lett. 99, 217402 (2007).
- Steffens et al. (2011) P. Steffens, O. Friedt, Y. Sidis, P. Link, J. Kulda, K. Schmalzl, S. Nakatsuji, and M. Braden, “Magnetic excitations in the metallic single-layer ruthenates Ca2-xSrxRuO4 studied by inelastic neutron scattering,” Phys. Rev. B 83, 054429 (2011).
- Friedt et al. (2001) O. Friedt, M. Braden, G. André, P. Adelmann, S. Nakatsuji, and Y. Maeno, “Structural and magnetic aspects of the metal-insulator transition in ,” Phys. Rev. B 63, 174432 (2001).
- Kriener et al. (2005) M. Kriener, P. Steffens, J. Baier, O. Schumann, T. Zabel, T. Lorenz, O. Friedt, R. Müller, A. Gukasov, P. G. Radaelli, P. Reutler, A. Revcolevschi, S. Nakatsuji, Y. Maeno, and M. Braden, “Structural Aspects of Metamagnetism in : Evidence for Field Tuning of Orbital Occupation,” Phys. Rev. Lett. 95, 267403 (2005).
- Ghosh et al. (2021) Sayak Ghosh, Arkady Shekhter, F. Jerzembeck, N. Kikugawa, Dmitry A. Sokolov, Manuel Brando, A.P. Mackenzie, Clifford W. Hicks, and B. J. Ramshaw, “Thermodynamic evidence for a two-component superconducting order parameter in Sr2RuO4,” Nat. Phys. 17, 199–204 (2021).
- Benhabib et al. (2021) Siham Benhabib, C. Lupien, I. Paul, L. Berges, M. Dion, Marc Nardone, Abdelaziz Zitouni, Z. Q. Mao, Y. Maeno, A. Georges, L. Taillefer, and C. Proust, “Ultrasound evidence for a two-component superconducting order parameter in Sr2RuO4,” Nat. Phys. 17, 194–198 (2021).
- Liechtenstein et al. (1996) A. I. Liechtenstein, O. Gunnarsson, O. K. Andersen, and R. M. Martin, “Quasiparticle bands and superconductivity in bilayer cuprates,” Phys. Rev. B 54, 12505–12508 (1996).
- Irkhin et al. (2002) V. Yu. Irkhin, A. A. Katanin, and M. I. Katsnelson, “Robustness of the Van Hove Scenario for High- Superconductors,” Phys. Rev. Lett. 89, 076401 (2002).
- Odashima et al. (2005) Satoru Odashima, Adolfo Avella, and Ferdinando Mancini, “High-order correlation effects in the two-dimensional Hubbard model,” Phys. Rev. B 72, 205121 (2005).
- Chen et al. (2011) K.-S. Chen, S. Pathak, S.-X. Yang, S.-Q. Su, D. Galanakis, K. Mikelsons, M. Jarrell, and J. Moreno, “Role of the van Hove singularity in the quantum criticality of the Hubbard model,” Phys. Rev. B 84, 245107 (2011).
- Schmitt (2010) Sebastian Schmitt, “Non-Fermi-liquid signatures in the Hubbard model due to van Hove singularities,” Phys. Rev. B 82, 155126 (2010).
- Blöchl (1994) P. E. Blöchl, “Projector augmented-wave method,” Phys. Rev. B 50, 17953–17979 (1994).
- Kresse and Joubert (1999) G. Kresse and D. Joubert, “From ultrasoft pseudopotentials to the projector augmented-wave method,” Phys. Rev. B 59, 1758–1775 (1999).
- Kresse and Furthmüller (1996) G. Kresse and J. Furthmüller, “Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set,” Comput. Mater. Sci. 6, 15–50 (1996).
- Kresse and Furthmüller (1996) G. Kresse and J. Furthmüller, “Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set,” Phys. Rev. B 54, 11169–11186 (1996).
- Perdew et al. (1996) John P. Perdew, Kieron Burke, and Matthias Ernzerhof, “Generalized Gradient Approximation Made Simple,” Phys. Rev. Lett. 77, 3865–3868 (1996).
- Marzari and Vanderbilt (1997) Nicola Marzari and David Vanderbilt, “Maximally localized generalized Wannier functions for composite energy bands,” Phys. Rev. B 56, 12847–12865 (1997).
- Marzari et al. (2012) Nicola Marzari, Arash A. Mostofi, Jonathan R. Yates, Ivo Souza, and David Vanderbilt, “Maximally localized Wannier functions: Theory and applications,” Rev. Mod. Phys. 84, 1419–1475 (2012).
- Mostofi et al. (2008) Arash A. Mostofi, Jonathan R. Yates, Young-Su Lee, Ivo Souza, David Vanderbilt, and Nicola Marzari, “wannier90: A tool for obtaining maximally-localised Wannier functions,” Comput. Phys. Commun. 178, 685–699 (2008).